# Constraints from Entanglement Wedge Nesting for Holography at a Finite Cutoff

### v1

@@ 5 Concluding Remarks
Our understanding of holography at a finite cutoff is poor and we hope that a careful study of properties of entanglement wedges in the presence of a cutoff can shed light on the properties of holographic duals to finite subregions. It would in particular be interesting to generalize our results to more complicated geometries and interval configurations, also in higher dimensions. Optimistically, this could provide a set of rules which determine under which conditions cutoff and boundary subregions can be treated as truly independent. It is interesting to note that violations of the conditions we obtained for EWN throughout this paper have always started in arbitrarily small bulk neighborhoods of the entangling surface on the cutoff. Perhaps this will always be the case if a non-trivial violation of EWN occurs, such that it could be sufficient to focus on those regions which would significantly simplify the analyses. Even if this is not generally the case one might at least formulate some concrete necessary conditions for EWN to hold again, this may provide insight as to how one can naturally identify independent subregions on the boundary.
@@ Acknowledgements.
I thank Niayesh Afshordi, Dominik Neuenfeld, Donald Marolf, Xi Dong, Samir Mathur, and Bin Guo for helpful discussions and comments as well as Dominik Neuenfeld for initial collaboration. The early stages of this work have also appeared in Krishan Saraswat’s Ph.D. thesis Saraswat:2023knb , which was supported by the University of Waterloo, Natural Sciences and Engineering Research Council of Canada (NSERC), and the Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities. From August 2023 - 2024 this work was supported by Clifford Johnson, the University of Southern California, and the US Department of Energy for support under grant DE-SC 0011687.
@@ A.2 Spacelike Separated Points From Line Segments AdS3
$\begin{split}&z^{(\pm)}_{\text{Null},i}(x,t)=\begin{cases}z^{(\pm)}_{\text{Null},\mathcal{R}_{R,i}}(x,t)&\frac{\partial_{x}z_{\sigma_{i}}}{|\partial_{x}z_{\sigma_{i}}|}(z_{\sigma_{i}}(x,t)-z^{(\pm)}_{\text{Null},\mathcal{R}_{R,i}}(x,t))\geq 0\\ z^{(\pm)}_{\text{Null},\mathcal{R}_{M,i}}(x,t)&\frac{\partial_{x}z_{\sigma_{i}}}{|\partial_{x}z_{\sigma_{i}}|}(z_{\sigma_{i}}(x,t)-z^{(\pm)}_{\text{Null},\mathcal{R}_{M,i}}(x,t))<0\\ \end{cases}\\ &z^{(\pm)}_{\text{Null},\mathcal{R}_{R,i}}(x,t)=b_{i}\cos\theta_{0}\pm\sqrt{(t-t_{b})^{2}-(x-b_{i}\sin\theta_{0})^{2}}\\ &z^{(\pm)}_{\text{Null},\mathcal{R}_{M,i}}(x,t)=\Theta[\pm(t_{i}(x)-t)]\sqrt{(t-t_{+,i})^{2}-(x-x_{+,i})^{2}}+\Theta[\pm(t-t_{i}(x))]\sqrt{(t-t_{-,i})^{2}-(x-x_{-,i})^{2}}\\ &z_{\sigma_{i}}(x,t)=\frac{2\Delta z_{i}}{\Delta t^{2}-\Delta x_{i}^{2}+\Delta z_{i}^{2}}\left(-x\Delta x_{i}+t\Delta t\right)+\frac{\Delta z_{i}\left[b_{i}^{2}-a_{i}^{2}+t_{a}^{2}-t_{b}^{2}\right]}{\Delta t^{2}-\Delta x_{i}^{2}+\Delta z_{i}^{2}}\\ &t_{i}(x)=t_{a}+\frac{\Delta t}{\Delta x_{i}}(x+a_{i}),\\ \end{split}$ (122)
@@ C.6 Condition for Connected Phase
$\begin{split}&\frac{\cos^{2}\tau_{\text{Bdry}}}{1-c_{t}^{2}}\frac{\sqrt{1-c_{t}^{2}\sin^{2}y_{\text{Br}}}-\sin y_{\text{Br}}\sqrt{1-c_{t}^{2}}}{\sqrt{1-c_{t}^{2}\sin^{2}y_{\text{Br}}}+\sin y_{\text{Br}}\sqrt{1-c_{t}^{2}}}\frac{\sinh\left(\frac{a^{\prime}+a}{z_{+}}\right)}{\sinh\left(\frac{a^{\prime}-a}{z_{+}}\right)}\sinh^{2}\left(\frac{a}{z_{+}}\right)\geq 1\\ &a^{\prime}=z_{+}\text{arccosh}\left[\frac{\cosh\left(\frac{a}{z_{+}}\right)}{\sqrt{1-\frac{\cos^{2}y_{\text{Br}}}{\cos^{2}\tau_{\text{Br}}}}}\right].\\ \end{split}$ (257)
In general we can see that as long as $a$ is sufficiently large, the inequality is satisfied, and therefore a connected phase is always guaranteed to exist as long as $a$ is sufficiently large. This makes the constraint we analyzed in the context of EWN relevant in the limit where $a\to\infty$ (i.e. the constraint we wrote in Eq. (84)).

### v2

@@ 5 Concluding Remarks
Our understanding of holography at a finite cutoff is poor and we hope that a careful study of properties of entanglement wedges in the presence of a cutoff can shed light on the properties of holographic duals to finite subregions. It would in particular be interesting to generalize our results to more complicated geometries and interval configurations, also in higher dimensions. Optimistically, this could provide a set of rules which determine under which conditions cutoff and boundary subregions can be treated as truly independent. It is interesting to note that violations of the conditions we obtained for EWN throughout this paper have always started in arbitrarily small bulk neighborhoods of the entangling surfaces. Perhaps this will always be the case if a non-trivial violation of EWN occurs, such that it could be sufficient to focus on those regions which would significantly simplify the analyses. Even if this is not generally the case one might at least formulate some concrete necessary conditions for EWN to hold again, this may provide insight as to how one can naturally identify independent subregions in cutoff holography. It may also be interesting to explore is how much our conclusions are affected through the particular model of cutoff surface we choose. Here we chose surfaces that are often employed in doubly holography setups and one might wonder if things may change for more general choices of cutoff surfaces [^23] It was pointed out in Omiya:2021olc that the ETW branes considered in double holography fails to satisfy the Gao-Wald theorem Gao:2000ga . We believe that our results can be regarded as a consequence of such violations of the Gao-Wald theorem. Based on this, we might speculate that more general cutoff surfaces also feature such violations of the Gao-Wald theorem and therefore our conclusions may carry over in a similar manner in more general contexts..
@@
Finally, another interesting avenue of research goes back to our discussion near the end of Section 1.1 of boundary conditions on the brane. In particular, when we impose Neumann boundary conditions on the brane we should think of the constraints that we obtain from EWN as constraints on brane world gravity. The rough procedure would be to understand how the location of the brane subregion, $B$, [^24] We fix the location of $A$ on a conformal boundary and location of $B$ is determined dynamically by understanding where the connected RT surface ends on the brane which should depend on the gravitational couplings on the brane. depends on various gravitational couplings in the brane-world theory of gravity. By precisely understanding this relation one could then straightforwardly translate the constraints we found here to constraints on gravity couplings. We leave such an investigation for later work.
@@ Acknowledgements.
I thank Niayesh Afshordi, Dominik Neuenfeld, Donald Marolf, Xi Dong, Vincent Chen, Samir Mathur, Bin Guo, Hao Geng, and Takato Mori for helpful discussions and comments as well as Dominik Neuenfeld for initial collaboration. The early stages of this work have also appeared in Krishan Saraswat’s Ph.D. thesis Saraswat:2023knb , which was supported by the University of Waterloo, Natural Sciences and Engineering Research Council of Canada (NSERC), and the Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities. From August 2023 - 2024 this work was supported by Clifford Johnson, the University of Southern California, and the US Department of Energy for support under grant DE-SC 0011687.
@@ A.2 Spacelike Separated Points From Line Segments AdS3
$\begin{split}&z^{(\pm)}_{\text{Null},i}(x,t)=\begin{cases}z^{(\pm)}_{\text{Null},\mathcal{R}_{R,i}}(x,t)\qquad\text{if}&\frac{\partial_{x}z_{\sigma_{i}}}{|\partial_{x}z_{\sigma_{i}}|}(z_{\sigma_{i}}(x,t)-z^{(\pm)}_{\text{Null},\mathcal{R}_{R,i}}(x,t))\geq 0\\ z^{(\pm)}_{\text{Null},\mathcal{R}_{M,i}}(x,t)\qquad\text{if}&\frac{\partial_{x}z_{\sigma_{i}}}{|\partial_{x}z_{\sigma_{i}}|}(z_{\sigma_{i}}(x,t)-z^{(\pm)}_{\text{Null},\mathcal{R}_{M,i}}(x,t))<0\\ \end{cases}\\ &z^{(\pm)}_{\text{Null},\mathcal{R}_{R,i}}(x,t)=b_{i}\cos\theta_{0}\pm\sqrt{(t-t_{b})^{2}-(x-b_{i}\sin\theta_{0})^{2}}\\ &z^{(\pm)}_{\text{Null},\mathcal{R}_{M,i}}(x,t)=\Theta[\pm(t_{i}(x)-t)]\sqrt{(t-t_{+,i})^{2}-(x-x_{+,i})^{2}}+\Theta[\pm(t-t_{i}(x))]\sqrt{(t-t_{-,i})^{2}-(x-x_{-,i})^{2}}\\ &z_{\sigma_{i}}(x,t)=\frac{2\Delta z_{i}}{\Delta t^{2}-\Delta x_{i}^{2}+\Delta z_{i}^{2}}\left(-x\Delta x_{i}+t\Delta t\right)+\frac{\Delta z_{i}\left[b_{i}^{2}-a_{i}^{2}+t_{a}^{2}-t_{b}^{2}\right]}{\Delta t^{2}-\Delta x_{i}^{2}+\Delta z_{i}^{2}}\\ &t_{i}(x)=t_{a}+\frac{\Delta t}{\Delta x_{i}}(x+a_{i}),\\ \end{split}$ (122)
@@ C.6 Condition for Connected Phase
$\begin{split}&\frac{\cos^{2}\tau_{\text{Bdry}}}{1-c_{t}^{2}}\left(\frac{\sqrt{1-c_{t}^{2}\sin^{2}y_{\text{Br}}}-\sin y_{\text{Br}}\sqrt{1-c_{t}^{2}}}{\sqrt{1-c_{t}^{2}\sin^{2}y_{\text{Br}}}+\sin y_{\text{Br}}\sqrt{1-c_{t}^{2}}}\right)\left(\frac{\sinh\left(\frac{a^{\prime}+a}{z_{+}}\right)}{\sinh\left(\frac{a^{\prime}-a}{z_{+}}\right)}\right)\sinh^{2}\left(\frac{a}{z_{+}}\right)\geq 1\\ &a^{\prime}=z_{+}\text{arccosh}\left[\frac{\cosh\left(\frac{a}{z_{+}}\right)}{\sqrt{1-\frac{\cos^{2}y_{\text{Br}}}{\cos^{2}\tau_{\text{Br}}}}}\right].\\ \end{split}$ (257)
Noting that $a^{\prime}\geq a$, by inspection, we can see that as long as $a$ is sufficiently large, the inequality is satisfied, and therefore a connected phase is always guaranteed to exist as long as $a$ is sufficiently large. This makes the constraint we analyzed in the context of EWN relevant in the limit where $a\to\infty$ (i.e. the constraint we wrote in Eq. (84)).